If 6M concentration means that there are 6 moles of HCl per liter of the solution and we know volume needed is 0.5 ml
what is the moles, molecular weight, and mass (g) of HCl?
Answer #1

1)

volume , V = 0.5 mL

= 5*10^-4 L

use:

number of mol,

n = Molarity * Volume

= 6*5*10^-4

= 3*10^-3 mol

Answer: 3*10^-3 mol

2)

Molar mass of HCl,

MM = 1*MM(H) + 1*MM(Cl)

= 1*1.008 + 1*35.45

= 36.458 g/mol

Answer: 36.458 g/mol

3)

use:

mass of HCl,

m = number of mol * molar mass

= 3*10^-3 mol * 36.46 g/mol

= 0.1094 g

Answer: 0.109 g